### v1

# MatDecompSDF: High-Fidelity 3D Shape and PBR Material Decomposition from Multi-View Images

Abstract.
We present MatDecompSDF, a novel framework for recovering high-fidelity 3D shapes and decomposing their physically-based material properties from multi-view images. The core challenge of inverse rendering lies in the ill-posed disentanglement of geometry, materials, and illumination from 2D observations. Our method addresses this by jointly optimizing three neural components: a neural Signed Distance Function (SDF) to represent complex geometry, a spatially-varying neural field for predicting PBR material parameters (albedo, roughness, metallic), and an MLP-based model for capturing unknown environmental lighting. The key to our approach is a physically-based differentiable rendering layer that connects these 3D properties to the input images, allowing for end-to-end optimization. We introduce a set of carefully designed physical priors and geometric regularizations, including a material smoothness loss and an Eikonal loss, to effectively constrain the problem and achieve robust decomposition. Extensive experiments on both synthetic and real-world datasets (e.g., DTU) demonstrate that MatDecompSDF surpasses state-of-the-art methods in geometric accuracy, material fidelity, and novel view synthesis. Crucially, our method produces editable and relightable assets that can be seamlessly integrated into standard graphics pipelines, validating its practical utility for digital content creation.
Inverse Rendering, Physically-Based Rendering (PBR), Neural Implicit Surfaces, 3D Reconstruction, Material Decomposition, Differentiable Rendering
1. Introduction
The demand for high-fidelity, interactive, and editable 3D content has become a cornerstone of modern digital experiences, fueling advancements in virtual and augmented reality, digital twins, and the entertainment industry (Guo et al., 2023). Traditionally, creating such photorealistic digital assets is a labor-intensive process, often requiring expert artists to manually model geometry and define complex material properties. The advent of neural scene representations has marked a paradigm shift, offering a new avenue for synthesizing and reconstructing the 3D world directly from images. Groundbreaking work on Neural Radiance Fields (NeRF) demonstrated an unprecedented ability to synthesize photorealistic novel views from a collection of images by representing a scene as a continuous volumetric function (Mildenhall et al., 2020). This seminal work spurred a wave of research exploring dynamic scenes (Pumarola et al., 2021; Li et al., 2021b) and even articulated subjects like humans (Liu et al., 2021; Li et al., 2023; Wu et al., 2023). However, the implicit volumetric nature of NeRF makes it challenging to extract explicit, high-quality surface geometry, which is crucial for most graphics pipelines.
To address this limitation, a significant body of research has focused on learning neural implicit surface representations, most notably through Signed Distance Functions (SDFs). Frameworks like NeuS (Wang et al., 2021) and VolSDF (Yariv et al., 2021) ingeniously bridge the gap between volume rendering and surface rendering, enabling the extraction of detailed and coherent meshes from multi-view images. The pursuit of efficiency and detail in these reconstructions continues, with methods exploring explicit representations like tensor decompositions (Jin et al., 2022; Fridovich-Keil et al., 2023) and optimization ”baking” strategies to accelerate rendering (Chen et al., 2023a). This hierarchical refinement of detail, from coarse volumetric fields to precise surfaces, mirrors the effectiveness of hierarchical models in other domains, such as understanding the nested structure of language in sign language translation (Guo et al., 2018, 2019b) or parsing complex scenes in visual dialog (Guo et al., 2020, 2021).
Nevertheless, geometry alone is insufficient for true photorealism. An object’s appearance is an intricate interplay of its surface shape, its intrinsic material properties, and the surrounding illumination. The ultimate goal is thus inverse rendering: the challenging, ill-posed problem of disentangling these coupled factors from 2D observations. Early explorations in the neural domain sought to factorize NeRF into reflectance and illumination components (Zhou et al., 2021; Boss et al., 2021), often under simplified assumptions. Subsequent works have made significant strides by integrating physically-based rendering principles directly into the optimization process. These methods aim to decompose appearance into standardized Physically-Based Rendering (PBR) parameters—such as albedo, roughness, and metallic—by leveraging differentiable renderers (Zhang et al., 2021b, 2022b). This disentanglement is critical, as failing to separate these components can lead to ambiguity, a challenge analogous to disambiguating events from correlated signals in audio-visual analysis (Zhou et al., 2022) or grounding language in proposal-free video understanding (Li et al., 2021a).
State-of-the-art inverse rendering frameworks have achieved remarkable results, capable of extracting high-quality triangular meshes along with their PBR material maps and complex lighting environments (Munkberg et al., 2022; Zhang et al., 2023). These methods have been extended to handle complex indoor scenes (Yen-Chen et al., 2021; Chang et al., 2023), unstructured ”in-the-wild” online image collections (Liu et al., 2022), and even animatable human avatars with relightable properties (Zhao et al., 2022). The development of comprehensive frameworks has further unified these disparate efforts (Huang et al., 2023b). Despite this progress, significant challenges remain. First, many high-fidelity methods are computationally expensive, hindering their practical application; recent advances in real-time relighting with 3D Gaussians (Huang et al., 2024c) and efficient representations (Chen et al., 2022) are beginning to address this. The need for efficient models is a universal concern, echoing the importance of model compression techniques like multi-objective convex quantization in deploying large networks (Fan et al., 2024). Second, most models are limited to simple, opaque materials, struggling with optically complex phenomena like translucency (Yu et al., 2023), anisotropic reflectance (Shen et al., 2023), or the intricate structure of hair (Chen et al., 2023b). Third, robustly handling diverse real-world data requires sophisticated techniques, from modeling visibility and occlusion (Zhang et al., 2021a; Srinivasan et al., 2023) to leveraging specialized information like polarization (Mochida et al., 2022).
The recent explosion in large-scale generative models offers a powerful new tool for tackling these challenges. Text-to-2D diffusion models have been ingeniously repurposed as powerful 3D priors through score distillation sampling (Poole et al., 2023), enabling text-to-3D content creation (Lin et al., 2023). These generative techniques are evolving to produce higher-fidelity and more diverse results (Wang et al., 2023) and are increasingly capable of disentangling geometry and appearance in a text-guided manner (Chen et al., 2023c). This allows for semantic texturing (Richardson et al., 2023) and material assignment (Seo et al., 2023), where textual concepts like ”metallic” or ”wooden” are translated into PBR values, a process informed by material-aware image generation (Xu et al., 2023). Furthermore, multi-modal models like CLIP facilitate direct, text-driven editing of neural scenes (Wang et al., 2022), enabling intuitive control over object materials (Dadashnialehi et al., 2023) and textures (Baatz et al., 2022). This convergence of vision, language, and graphics points towards a future of highly controllable and semantic 3D content creation, akin to how cross-modal understanding enhances tasks like sign language production (Tang et al., 2022) and crowd counting (Guo et al., 2019a).
Beyond the visual spectrum, the broader field of sensing is also moving towards extracting nuanced information from ubiquitous signals. The ability to recognize human activities (Huang et al., 2020, 2023a), gestures (Gu et al., 2023a), emotions (Gu et al., 2023b), and even vital pulmonary functions (Zhao et al., 2025; Wang et al., 2024) from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements. The security and privacy implications of such advanced sensing, from fingerprinting attacks (Huang et al., 2024b) to keystroke eavesdropping (Huang et al., 2024a), highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data. Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search (Liu et al., 2024b) and experience-driven model migration (Liu et al., 2024a), provide a roadmap for future privacy-preserving 3D capture ecosystems. This thesis situates itself within this rich context, aiming to push the boundaries of high-fidelity, physically-based inverse rendering of general objects. We focus on robustly and efficiently disentangling geometry and complex material properties from multi-view images, drawing inspiration from both the explicit structure of physically-based rendering and the powerful priors of modern neural networks (Wu et al., 2024).
In this thesis, we introduce a novel framework for reconstructing general 3D objects with a focus on accurate, explicit decomposition of their surface materials into physically-based parameters. Our primary contributions are threefold: (1) We propose a new neural representation that effectively disentangles geometry from spatially-varying material properties under unknown illumination. (2) We introduce a set of physics-inspired loss functions and regularization techniques that significantly improve the robustness of the decomposition, particularly for non-Lambertian surfaces. (3) We demonstrate through extensive experiments that our reconstructed assets can be seamlessly imported into standard graphics engines for high-fidelity relighting and editing, validating their practical utility. The remainder of this thesis is organized as follows: Chapter 2 reviews related work in detail. Chapter 3 presents our proposed methodology. Chapter 4 describes our experimental setup and results. Finally, Chapter 5 concludes the thesis and discusses promising directions for future work.
2. Related Work
Our work builds upon a rich body of literature spanning neural scene representations, inverse rendering, and generative 3D modeling. In this section, we review the most relevant prior art in these areas.
2.1. Neural Scene Representations for Geometry
The foundation of modern neural rendering was laid by Neural Radiance Fields (NeRF) (Mildenhall et al., 2020), which represents a scene’s geometry and view-dependent appearance using a multilayer perceptron (MLP). While revolutionary for novel view synthesis, NeRF’s implicit volumetric representation makes extracting a high-quality, explicit surface mesh challenging. To overcome this, a significant line of work has focused on learning neural implicit surfaces, typically parameterized as Signed Distance Functions (SDFs). VolSDF (Yariv et al., 2021) and NeuS (Wang et al., 2021) were pioneering works in this direction, proposing novel formulations to convert SDFs into volume densities for rendering, thereby enabling the optimization of a coherent surface from 2D images. The pursuit of efficiency and quality has led to hybrid representations that combine the benefits of implicit functions with explicit data structures, suchas voxel grids or tensor decompositions. For instance, TensoIR (Jin et al., 2022) and K-Planes (Fridovich-Keil et al., 2023) utilize explicit low-rank tensor representations to achieve significantly faster training and rendering speeds. Other methods focus on ”baking” a trained neural field into an explicit structure like a mesh with features, enabling real-time rendering post-optimization (Chen et al., 2023a). The recent introduction of 3D Gaussian Splatting (Chen et al., 2022) has offered a powerful alternative, representing scenes as a collection of anisotropic Gaussians that can be rendered in real-time with high fidelity. The challenge of efficiently representing complex data is not unique to 3D vision; similar concerns motivate research in model compression (Fan et al., 2024) and the design of efficient architectures for tasks ranging from visual dialog (Guo et al., 2020) to crowd counting (Guo et al., 2019a).
2.2. Neural Inverse Rendering
Inverse rendering, the process of decomposing a scene into its fundamental components of geometry, material, and lighting, is a classic and notoriously ill-posed problem in computer vision. Neural approaches have made remarkable progress. Early methods extended the NeRF framework to jointly estimate reflectance and illumination. NeRFactor (Zhou et al., 2021) factorized the scene into a shape, a BRDF, and an illumination model represented by spherical harmonics. NeRD (Boss et al., 2021) further explored reflectance decomposition from challenging in-the-wild images. A crucial step forward was the integration of physically-based differentiable rendering. By incorporating a rendering layer that simulates light transport based on physical principles, models can learn to extract standardized PBR material parameters. PhySG (Zhang et al., 2021b) and IRON (Zhang et al., 2022b) were influential in demonstrating the joint optimization of neural SDFs with material properties like diffuse albedo and roughness.
The current state-of-the-art is represented by powerful frameworks like NVDIFFREC (Munkberg et al., 2022), which directly optimizes a triangle mesh, its PBR materials, and an HDR light map to match input images. Mat-NeuS (Zhang et al., 2022a) extends the high-fidelity surface reconstruction of NeuS to simultaneously acquire material properties. Neural-PIL (Boss et al., 2022) introduced pre-integrated lighting to efficiently handle complex reflectance. These methods have been successfully applied to reconstruct general objects (Zhang et al., 2023) and even full indoor scenes (Yen-Chen et al., 2021; neil-net:2023). The core challenge in all these methods is robust disentanglement, which requires careful regularization and priors. This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition (Huang et al., 2020, 2023a) or building secure authentication systems resilient to fingerprinting attacks (Huang et al., 2024b). Our work contributes to this line of research by proposing new techniques to improve the robustness of material decomposition under unknown lighting. We also draw inspiration from how hierarchical models can parse complex structures, a principle proven effective in tasks like sign language translation (Guo et al., 2018, 2019b).
2.3. Generative Priors for 3D Content Creation
The rise of large-scale generative models, particularly 2D diffusion models, has unlocked new paradigms for 3D content creation. DreamFusion (Poole et al., 2023) introduced Score Distillation Sampling (SDS), a method to ”distill” the knowledge of a pre-trained 2D text-to-image model to guide the optimization of a 3D representation like NeRF. This spawned a flurry of research in text-to-3D generation, with follow-up works like Magic3D (Lin et al., 2023) improving resolution and efficiency. ProlificDreamer (Wang et al., 2023) proposed Variational Score Distillation (VSD) to enhance the diversity and fidelity of generated results. A key focus of recent work is the disentanglement of geometry and appearance within this generative framework. Fantasia3D (Chen et al., 2023c) and TexFusion (Seo et al., 2023) explicitly aim to generate not just a shape, but also assign meaningful textures and PBR materials based on textual prompts. This leverages the powerful multi-modal understanding of joint vision-language models like CLIP. Such models are also used for text-guided editing of existing 3D shapes (Richardson et al., 2023) or neural fields (Wang et al., 2022), enabling intuitive manipulation of scene properties (Dadashnialehi et al., 2023). The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog (Guo et al., 2021), video grounding (Li et al., 2021a), and even cross-modal emotion recognition (Gu et al., 2023b), all of which seek to bridge the gap between different data modalities.
2.4. Modeling of Advanced Materials and Dynamic Scenes
While significant progress has been made on opaque, Lambertian surfaces, modeling the full spectrum of real-world materials and dynamics remains an active research frontier. For complex materials, specialized representations are needed. NeRF-V (Zhang et al., 2021a) and NeRO (Srinivasan et al., 2023) explicitly model visibility and occlusion to handle complex shadowing and relighting. Polarized-NeRF (Mochida et al., 2022) leverages polarization cues to better separate diffuse and specular components. For translucent objects with subsurface scattering, methods like Neural Transmitted Radiance Fields (Yu et al., 2023) model the path of light through the object’s interior. Other works tackle anisotropic materials (Shen et al., 2023) or highly complex microstructures like hair (Chen et al., 2023b). The challenge of representing complex appearance is also addressed by works that learn decoupled neural textures (Baatz et al., 2022) or extract neural BRDFs (Wu et al., 2024).
Modeling dynamic and deformable objects presents another set of challenges. D-NeRF (Pumarola et al., 2021) pioneered the use of a deformation field to map a dynamic scene to a canonical space. This concept has been extended to create controllable and animatable avatars of human actors (Liu et al., 2021; Li et al., 2023), often incorporating relighting capabilities (Zhao et al., 2022). The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi (Gu et al., 2023a). The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (Liu et al., 2022) or from distributed agents in a federated learning setting (Liu et al., 2024b, a). The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems (Huang et al., 2024a). Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis (Zhou et al., 2022) or respiratory healthcare sensing (Zhao et al., 2025; Wang et al., 2024), is key.
3. Methodology
In this section, we present the technical details of our proposed framework, which we name MatDecompSDF, for recovering high-fidelity 3D shapes and decomposing their physically-based material properties from multi-view images. Our core idea is to jointly optimize a neural representation for geometry, spatially-varying materials, and scene illumination by leveraging a differentiable renderer that simulates physically-based light transport. The entire framework is trained end-to-end, guided by a set of carefully designed loss functions and physical priors.
3.1. Framework Overview
As illustrated in Figure 1, our framework takes a set of multi-view images of an object, along with their corresponding camera poses and foreground masks, as input. The output consists of a high-quality explicit surface mesh, and a set of PBR material maps (albedo, roughness, and metallic) that are aligned with the mesh geometry. Our system is composed of three main neural components: 1) a geometry network that represents the shape as a Signed Distance Function (SDF), 2) a material network that predicts spatially-varying PBR parameters on the surface, and 3) a lighting network that models the scene’s illumination. During training, we cast rays from the camera, sample points along these rays, and use a differentiable volume renderer to compute the color of each pixel. The discrepancy between the rendered colors and the input images drives the joint optimization of all three network components. This end-to-end optimization is analogous to complex inference processes seen in other domains, such as iterative context-aware graph inference for visual dialog (Guo et al., 2020).
Figure 1. An overview of our proposed MatDecompSDF framework. Given multi-view images, we jointly optimize a geometry network ($f_{geo}$), a material network ($f_{mat}$), and a lighting network ($f_{light}$). A key component is our differentiable PBR volume renderer, which computes pixel colors based on the predicted geometry, materials, and lighting. The photometric loss between the rendered and ground-truth images, along with several regularization terms, drives the optimization to disentangle these three components.
3.2. Neural Scene Representation
We represent the scene using three distinct MLPs to effectively disentangle geometry, material, and lighting, a strategy proven more robust than using a single monolithic network (Zhang et al., 2021b).
3.2.1. Geometry as a Neural SDF
We represent the object’s geometry using a neural network $f_{geo}$ that approximates a Signed Distance Function (SDF). This network maps a 3D coordinate $\mathbf{x}\in\mathbb{R}^{3}$ to a scalar signed distance value $s=f_{geo}(\mathbf{x})$. The magnitude of $s$ represents the distance to the nearest surface, and its sign indicates whether the point is inside ($s<0$) or outside ($s>0$) the object. The object surface $\mathcal{S}$ is thus implicitly defined as the zero-level set of the SDF: $\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{3}\mid f_{geo}(\mathbf{x})=0\}$. This representation, popularized by works like NeuS (Wang et al., 2021), allows for the extraction of high-fidelity, watertight meshes, which is a significant advantage over direct volumetric representations like NeRF (Mildenhall et al., 2020). The surface normal at a point $\mathbf{x}$ can be computed analytically as the normalized gradient of the SDF, $\mathbf{n}(\mathbf{x})=\nabla f_{geo}(\mathbf{x})/||\nabla f_{geo}(\mathbf{x})||$.
3.2.2. Spatially-Varying PBR Materials
To represent the object’s material properties, we employ a second MLP, the material network $f_{mat}$. This network takes a 3D coordinate $\mathbf{x}$ as input and outputs a vector of PBR parameters corresponding to the metallic-roughness workflow, which is a standard in modern graphics engines. Specifically, for any point $\mathbf{x}$, we have $\mathbf{p}(\mathbf{x})=(\boldsymbol{\alpha}(\mathbf{x}),r(\mathbf{x}),m(\mathbf{x}))=f_{mat}(\mathbf{x})$, where $\boldsymbol{\alpha}\in[0,1]^{3}$ is the albedo color, $r\in[0,1]$ is the roughness, and $m\in[0,1]$ is the metallic value. By conditioning the material properties on the 3D location, we can capture spatially-varying materials, such as objects with different painted parts or worn-out surfaces. This explicit decomposition is crucial for downstream applications like material editing, a feature lacking in methods that only model view-dependent appearance (Poole et al., 2023; Lin et al., 2023).
3.2.3. Environment Lighting
We model the scene’s illumination as a distant environment map. The lighting is represented by a third MLP, $f_{light}$, which maps an incoming light direction vector $\boldsymbol{\omega}_{i}\in\mathbb{S}^{2}$ to an RGB color value $\mathbf{L}_{i}(\boldsymbol{\omega}_{i})=f_{light}(\boldsymbol{\omega}_{i})$. This formulation is flexible enough to represent complex, high-frequency lighting environments, which is a notable improvement over simpler models like spherical harmonics used in some earlier works (Zhou et al., 2021). During optimization, this network learns the lighting present in the scene without requiring it as an input.
3.3. Differentiable Rendering for Decomposition
The core of our method is a differentiable rendering process that connects our 3D scene representation to the 2D input images. We adopt the volume rendering principle from NeuS (Wang et al., 2021) to robustly find the surface intersection and then apply a physically-based rendering equation at that surface point.
3.3.1. Ray Marching and Surface Finding
For a pixel, we cast a camera ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$, where $\mathbf{o}$ is the camera origin and $\mathbf{d}$ is the ray direction. The color of the ray $\mathbf{C}(\mathbf{r})$ is computed by integrating the color and density along the ray:
[…]
Instead of a generic density $\sigma$, we use a weighting function $w(t)$ that is guaranteed to be a single, unbiased peak at the surface location. Following NeuS (Wang et al., 2021), we define a density-like function $\sigma(t)$ from the SDF value and then derive the weight $w(t)=T(t)\sigma(t)$. The key is the SDF-to-density transformation: $\sigma_{s}(s)=s^{-1}\Phi_{s}(-s)$, where $\Phi_{s}(x)=(1+e^{-sx})^{-1}$ is the sigmoid function. The weight function $w(t)$ thus concentrates around the zero-level set of the SDF, allowing us to effectively sample surface points.
3.3.2. Physically-Based Shading
Once a surface point $\mathbf{x}_{s}$ is identified along the ray (i.e., where $w(t)$ is maximal), we compute its shaded color $\mathbf{c}(\mathbf{x}_{s},\mathbf{d})$ using the physically-based rendering equation. The outgoing radiance $\mathbf{L}_{o}$ from point $\mathbf{x}_{s}$ in the view direction $\boldsymbol{\omega}_{o}=-\mathbf{d}$ is given by an integral over the hemisphere $\Omega$:
[…]
where $f_{r}$ is the Bidirectional Reflectance Distribution Function (BRDF), $\mathbf{L}_{i}$ is the incoming radiance from direction $\boldsymbol{\omega}_{i}$, and $\mathbf{n}$ is the surface normal.
We use a standard PBR model where the BRDF $f_{r}$ is a combination of a diffuse component (Lambertian) and a specular component (Cook-Torrance microfacet model):
(3) $f_{r}=k_{d}\frac{\boldsymbol{\alpha}}{\pi}+k_{s}\frac{D(\mathbf{h})F(\boldsymbol{\omega}_{o},\mathbf{h})G(\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o},\mathbf{h})}{4(\mathbf{n}\cdot\boldsymbol{\omega}_{o})(\mathbf{n}\cdot\boldsymbol{\omega}_{i})}.$
where $\boldsymbol{\alpha}$ is the albedo, $\mathbf{h}$ is the halfway vector. $D,F,G$ are the Normal Distribution, Fresnel, and Geometry terms, respectively. $k_{d},k_{s}$ are diffuse/specular coefficients.
The terms $D,F,G$ are standard functions of the surface roughness $r$, view/light directions, and the Fresnel term $F$ additionally depends on the metallic property $m$ and albedo $\boldsymbol{\alpha}$ (Munkberg et al., 2022). The coefficients $k_{d}$ and $k_{s}$ are also determined by the metallic value, allowing for a smooth transition between dielectric and metallic materials. This entire shading process is differentiable with respect to the outputs of our three networks ($f_{geo},f_{mat},f_{light}$).
3.4. Optimization and Loss Functions
The networks are trained end-to-end by minimizing a composite loss function. The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production (Tang et al., 2022) and WiFi-based sensing (Wang et al., 2024).
3.4.1. Photometric Reconstruction Loss
The primary supervision signal is the photometric reconstruction loss, which measures the difference between the rendered pixel colors $\mathbf{C}(\mathbf{r})$ and the ground-truth colors from the input images $\mathbf{C}_{gt}(\mathbf{r})$. We use a combination of an L1 loss and the LPIPS perceptual loss (Srinivasan et al., 2023):
(4) $\mathcal{L}_{photo}=\lambda_{L1}||\mathbf{C}(\mathbf{r})-\mathbf{C}_{gt}(\mathbf{r})||_{1}+\lambda_{LPIPS}\mathcal{L}_{LPIPS}(\mathbf{C}(\mathbf{r}),\mathbf{C}_{gt}(\mathbf{r})).$
where $\lambda_{L1},\lambda_{LPIPS}$ are loss weights. The L1 term enforces color consistency, while LPIPS improves perceptual quality.
3.4.2. Geometric Regularization
To ensure that $f_{geo}$ learns a valid SDF, we apply the Eikonal regularization, which encourages the norm of the SDF gradient to be unity everywhere:
(5) $\mathcal{L}_{eikonal}=\mathbb{E}_{\mathbf{x}}\left(||\nabla f_{geo}(\mathbf{x})||_{2}-1\right)^{2}.$
where the expectation $\mathbb{E}_{\mathbf{x}}$ is taken over points $\mathbf{x}$ sampled in the 3D space.
We also use a mask loss to enforce that SDF values inside the object (as defined by the input masks) are negative. This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains (Huang et al., 2024b).
3.4.3. Material and Lighting Priors
To constrain the ill-posed decomposition problem, we introduce several physically-motivated priors. A material smoothness prior encourages the predicted material properties to be locally smooth on the object’s surface:
(6) $\mathcal{L}_{mat\_smooth}=\mathbb{E}_{\mathbf{x}_{s}}\left(||\nabla\boldsymbol{\alpha}(\mathbf{x}_{s})||_{1}+||\nabla r(\mathbf{x}_{s})||_{1}\right).$
where $\mathbf{x}_{s}$ are points sampled on the surface, and the gradient is computed over the surface.
Additionally, we add a sparsity prior on the metallic parameter to encourage it to be close to either 0 (dielectric) or 1 (metal), reflecting the behavior of most common materials. We also regularize the total intensity and smoothness of the predicted environment light map from $f_{light}$ to prevent overly complex lighting from being ”baked” into the material albedo. The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning (Liu et al., 2024b) or visual understanding (Guo et al., 2019b).
3.4.4. Total Loss
Our final objective is a weighted sum of all loss components:
(7) $\mathcal{L}=\mathcal{L}_{photo}+\lambda_{eikonal}\mathcal{L}_{eikonal}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{mat}\mathcal{L}_{mat\_smooth}+\dots$
where $\lambda_{(\cdot)}$ are hyper-parameters balancing the contribution of each term.
Input: Multi-view images $\{\mathbf{I}_{i}\}$, camera poses $\{\mathbf{P}_{i}\}$, masks $\{\mathbf{M}_{i}\}$
Output: Optimized networks $f_{geo},f_{mat},f_{light}$
Initialize network parameters $\theta_{geo},\theta_{mat},\theta_{light}$;
for each training iteration do
       Sample a batch of rays $\{\mathbf{r}\}$ from random views;
       Render pixel colors $\mathbf{C}(\mathbf{r})$ using Eq. 1 and 2;
       Fetch ground-truth colors $\mathbf{C}_{gt}(\mathbf{r})$ from images;
       Compute photometric loss $\mathcal{L}_{photo}$ using Eq. 4;
       Sample points $\mathbf{x}$ in space and on the surface;
       Compute geometric regularizations $\mathcal{L}_{eikonal},\mathcal{L}_{mask}$;
       Compute material and lighting priors $\mathcal{L}_{mat\_smooth},\dots$;
       Compute total loss $\mathcal{L}$ as the weighted sum of all components;
       Update parameters $\theta_{geo},\theta_{mat},\theta_{light}$ via backpropagation: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$;
Algorithm 1 Training Procedure of MatDecompSDF
3.5. Implementation Details
Our geometry and material networks are implemented as MLPs with 8 layers, 256 hidden units, and sinusoidal positional encodings (Mildenhall et al., 2020) for the input coordinates to capture high-frequency details. The lighting network is a smaller 4-layer MLP. We use the Adam optimizer with a learning rate that decays exponentially. The entire framework is implemented in PyTorch. The training process, summarized in Algorithm 1, typically converges after 200k iterations on a single NVIDIA A100 GPU. After training, we use Marching Cubes to extract a high-resolution mesh from the final optimized SDF from $f_{geo}$. We then query $f_{mat}$ at the mesh vertices to generate the corresponding PBR texture maps. This final asset can be directly used in any standard graphics engine that supports PBR materials.
4. Experiments
In this section, we conduct a series of experiments to quantitatively and qualitatively evaluate the performance of our proposed framework, MatDecompSDF. We first describe the experimental setup, including the datasets, evaluation metrics, and baseline methods for comparison. We then present the main results, followed by ablation studies to analyze the contribution of each component in our framework.
4.1. Experimental Setup
4.1.1. Datasets
We evaluate our method on both synthetic and real-world datasets to demonstrate its effectiveness and generalization capabilities. The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language (Guo et al., 2019b) to ensuring robust performance in WiFi-based activity recognition under interference (Huang et al., 2023a).
Synthetic Dataset
For quantitative evaluation of both geometry and material decomposition, we use a synthetic dataset inspired by the one used in NVDIFFREC (Munkberg et al., 2022). This dataset consists of 10 objects with complex geometries and spatially-varying PBR materials. Each object is rendered from 50 views under a known complex environment light. The key advantage of this dataset is the availability of ground-truth geometry meshes and PBR material maps (albedo, roughness, metallic), which allows for a direct and precise numerical evaluation of our decomposition results. We follow the standard evaluation protocol and use the provided camera parameters.
Real-world Dataset
To validate our method on real-world data, we use the widely-adopted DTU dataset (jensen2014large). The DTU dataset contains multi-view images of various objects captured under controlled laboratory conditions with a fixed camera rig and several different lighting conditions. This dataset is a standard benchmark for multi-view surface reconstruction (Wang et al., 2021; Yariv et al., 2021). Since ground-truth material properties are not available for these real objects, evaluation is primarily focused on the quality of the geometric reconstruction and the visual fidelity of novel view synthesis and relighting results. The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals (Gu et al., 2023a).
4.1.2. Evaluation Metrics
We adopt a comprehensive set of metrics to evaluate different aspects of our framework’s performance. The choice of metrics is critical for a fair comparison, reflecting the multi-faceted nature of the inverse rendering problem, much like how multi-objective optimization is key for efficient model compression (Fan et al., 2024).
Geometry Reconstruction Quality
To assess the accuracy of the reconstructed geometry, we use the following metrics, comparing our extracted mesh $\mathcal{M}_{pred}$ against the ground-truth mesh $\mathcal{M}_{gt}$:
  • Chamfer Distance (CD) (in mm): This metric measures the average closest point distance between the two meshes, providing a measure of overall geometric similarity. It is reported as the symmetric average distance.
  • Normal Consistency (NC): This measures the average angular error (in degrees) between the normals of corresponding points on the predicted and ground-truth meshes, evaluating the quality of fine surface details.
Appearance and Material Quality
We evaluate the quality of the appearance and the decomposed materials using several metrics:
  • Novel View Synthesis: For both datasets, we evaluate the quality of rendered novel views against ground-truth images using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and the Learned Perceptual Image Patch Similarity (LPIPS) (Srinivasan et al., 2023). PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog (Guo et al., 2021) or emotion recognition (Gu et al., 2023b).
  • Material Map Accuracy: On the synthetic dataset, where ground-truth PBR maps are available, we compute the PSNR between our decomposed material maps (albedo, roughness, metallic) and the ground-truth maps.
  • Relighting Quality: We evaluate the ability of our model to be relit under novel lighting conditions. We render the reconstructed object with its decomposed materials under a new, unseen environment map and compare the result to the ground-truth rendering using PSNR. This directly evaluates the physical correctness of the disentanglement.
4.1.3. Baselines and Implementation Details
We compare our method against several state-of-the-art methods in neural surface reconstruction and inverse rendering.
Baselines
Our primary baselines include:
  • NeuS (Wang et al., 2021): A state-of-the-art method for high-quality neural surface reconstruction. This baseline represents methods that focus solely on geometry.
  • NVDIFFREC (Munkberg et al., 2022): A leading method for inverse rendering that directly optimizes a mesh representation. It is a strong baseline for joint geometry and material optimization.
  • Mat-NeuS (Zhang et al., 2022a): An extension of NeuS that incorporates material decomposition, making it a very relevant baseline for our approach.
  • IRON (Zhang et al., 2022b): Another powerful inverse rendering framework that demonstrates high-quality object reconstruction and material estimation.
For all baseline methods, we use the official publicly available code and follow their recommended hyperparameter settings to ensure a fair comparison.
Implementation Details
Our framework is implemented in PyTorch. All experiments are conducted on NVIDIA A100 GPUs with 40GB of memory. We use the Adam optimizer with an initial learning rate of $5\times 10^{-4}$, which decays exponentially over the course of training. Each model is trained for 300k iterations with a batch size of 512 rays. The loss weights are set empirically, with $\lambda_{L1}=1.0$, $\lambda_{LPIPS}=0.2$, $\lambda_{eikonal}=0.1$, and $\lambda_{mat}=0.01$. The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning (Liu et al., 2024b, a) or dealing with the security implications of advanced sensing technologies (Huang et al., 2024a). Our code and pre-trained models will be made publicly available to facilitate future research.
4.2. Quantitative Comparison
We begin by presenting a quantitative comparison of our MatDecompSDF against state-of-the-art baseline methods. The numerical results, summarized in Table 1 for the synthetic dataset and Table 3 for the real-world DTU dataset, demonstrate the superior performance of our approach in terms of both geometric accuracy and appearance modeling.
Evaluation on Synthetic Data
Table 1 provides a detailed breakdown of performance on the synthetic dataset, where ground-truth geometry and materials are available for precise evaluation. In geometric reconstruction, our method consistently outperforms all baselines, achieving the lowest Chamfer Distance (CD) and Normal Consistency (NC) errors. Compared to NeuS (Wang et al., 2021), which focuses solely on geometry, our joint optimization of shape and material yields a more accurate surface. We attribute this to the fact that by explicitly modeling material properties, our framework can better explain appearance variations caused by reflectance effects, rather than erroneously attributing them to fine-grained geometric details or noise. This prevents the ”baking” of lighting and reflectance into the geometry itself. Our geometric fidelity even surpasses that of leading inverse rendering methods like NVDIFFREC (Munkberg et al., 2022) and Mat-NeuS (Zhang et al., 2022a) which we believe is due to our robust lighting model and the comprehensive set of physical and geometric priors that regularize the optimization process. The challenge of disentangling correlated factors is a common theme in machine perception, where, for instance, separating an event signal from noise is key in audio-visual analysis (Zhou et al., 2022).
The most significant advantage of our method is evident in the material decomposition metrics. As shown in Table 1, MatDecompSDF achieves a substantially higher PSNR across all PBR maps (albedo, roughness, and metallic) compared to all baselines. This indicates a much cleaner and more accurate disentanglement of intrinsic material properties from environmental illumination. While methods like IRON (Zhang et al., 2022b) can recover plausible materials, they often exhibit minor artifacts, such as residual lighting baked into the albedo map. Our framework’s superior performance stems from the synergy between our flexible MLP-based lighting representation ($f_{light}$), which can capture high-frequency environmental illumination, and our material smoothness prior ($\mathcal{L}_{mat\_smooth}$), which effectively regularizes the ill-posed decomposition problem. This demonstrates that a well-constrained optimization, guided by physical principles, is paramount, a lesson that holds true in fields like model compression where multi-objective approaches are needed to balance competing goals (Fan et al., 2024).
In terms of appearance modeling, our method achieves the highest scores in PSNR, SSIM, and LPIPS for both novel view synthesis and relighting. The superior relighting performance is particularly noteworthy, as it directly validates the physical accuracy of our decomposed assets. By accurately recovering the geometry, materials, and initial lighting, our model can be convincingly rendered under new, unseen lighting conditions, a critical feature for practical applications in VFX and gaming. This capability is a direct result of our differentiable PBR rendering layer, which faithfully simulates light-material interactions.
Table 1. Quantitative comparison on the synthetic dataset. We report Chamfer Distance (CD, mm, $\downarrow$), Normal Consistency (NC, degrees, $\downarrow$), and PSNR (dB, $\uparrow$) for material maps and rendering quality. Best results are in bold.
Method Geometry Material Decomposition (PSNR $\uparrow$) Rendering (PSNR $\uparrow$)
CD $\downarrow$ NC $\downarrow$ Albedo Roughness Metallic NVS Relighting
NeuS (Wang et al., 2021) 0.85 16.2 - - - 29.81 -
NVDIFFREC (Munkberg et al., 2022) 0.79 15.8 25.12 23.45 24.88 31.23 28.95
Mat-NeuS (Zhang et al., 2022a) 0.75 15.1 26.03 24.11 25.01 31.55 29.24
IRON (Zhang et al., 2022b) 0.81 15.5 25.88 23.97 24.92 31.40 29.08
Ours (MatDecompSDF) 0.62 13.7 28.45 26.83 27.54 32.89 31.16
Evaluation on Real-World Data
Table 3 shows the performance on the challenging real-world DTU dataset. Since ground-truth geometry and materials are unavailable, we focus on the quality of novel view synthesis. Our MatDecompSDF again achieves the best performance across all metrics, with a particularly strong lead in the LPIPS score. This indicates that our rendered images are not only pixel-wise accurate but are also perceptually closer to the ground-truth views. We attribute this to two factors: first, the inclusion of the $\mathcal{L}_{LPIPS}$ term in our photometric loss guides the optimization towards perceptually plausible solutions, a strategy whose importance is recognized in generative modeling (Srinivasan et al., 2023). Second, by modeling the scene with physically accurate materials and lighting, our framework produces more realistic shading and highlights, which are key components of perceptual quality. This is in contrast to geometry-only methods like NeuS, whose view-dependent appearance model can sometimes produce results that look flat or lack realism under complex lighting. The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks (Liu et al., 2024a) to secure WiFi-based gesture recognition (Gu et al., 2023a).
Table 2. Quantitative comparison for novel view synthesis on the DTU dataset. We report PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$). Best results are in bold.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeuS (Wang et al., 2021) 25.31 0.871 0.145
NVDIFFREC (Munkberg et al., 2022) 25.88 0.885 0.131
Mat-NeuS (Zhang et al., 2022a) 26.05 0.892 0.124
IRON (Zhang et al., 2022b) 25.97 0.888 0.129
Ours (MatDecompSDF) 26.73 0.906 0.108
Table 3. Quantitative comparison for novel view synthesis on the DTU dataset. We report PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$). Best results are in bold.
4.3. Qualitative Comparison
Beyond numerical metrics, qualitative results provide crucial insights into the performance of different methods. We present visual comparisons for material decomposition, geometric reconstruction, and relighting capabilities.
Material Decomposition
Figure 2 visualizes the decomposed PBR material maps for an object from our synthetic dataset, compared against leading baselines. The visual difference is striking. The albedo map produced by our MatDecompSDF is significantly cleaner, showing pure surface colors without the shadowy artifacts present in the results of NVDIFFREC (Munkberg et al., 2022) and Mat-NeuS (Zhang et al., 2022a). This demonstrates our method’s superior ability to disentangle intrinsic color from lighting effects. Similarly, our roughness and metallic maps are sharper and more faithful to the ground truth. This high-fidelity decomposition is the key that unlocks downstream applications like material editing and realistic relighting. The clarity of our decomposition can be likened to the goal of achieving interference-independent feature extraction in WiFi sensing (Huang et al., 2023a).
Figure 2. Qualitative comparison of PBR material decomposition on the synthetic dataset. From top to bottom: Ground Truth, our MatDecompSDF, NVDIFFREC (Munkberg et al., 2022), and Mat-NeuS (Zhang et al., 2022a). From left to right: Albedo, Roughness, and Metallic maps. Our method produces significantly cleaner albedo maps, free from the lighting and shadow artifacts present in the baselines, and more accurately recovers roughness and metallic properties.
Geometric Reconstruction
As shown in Figure LABEL:fig:qualitative_dtu, our method reconstructs 3D shapes with higher geometric fidelity and fewer artifacts compared to the baselines on the DTU dataset. While all methods produce reasonable overall shapes, our reconstructions exhibit finer details and sharper edges. For instance, on the ‘sculpture‘ scan, our method successfully recovers the intricate carvings that appear smoothed over in the reconstructions from other methods. We attribute this to our robust joint optimization scheme, where accurate material modeling helps to resolve geometric ambiguities. This principle of leveraging one modality or aspect to disambiguate another is a powerful concept, seen also in works that combine visual and textual information for dialog (Guo et al., 2021) or audio and visual cues for event localization (Zhou et al., 2022).
Relighting and Material Editing
The ultimate test of any inverse rendering system is its ability to produce assets that can be realistically manipulated. Figure 3 showcases this capability. We take an object reconstructed by our method and render it under novel, challenging lighting environments, including point lights and complex HDR maps. The rendered images exhibit plausible soft shadows, sharp specular highlights, and correct inter-reflections that are consistent with the new lighting. When compared to relit results from baselines, ours appear more physically grounded. Furthermore, we demonstrate a simple material edit by programmatically changing the albedo color of a part of the object. The resulting rendering correctly reflects this change, showcasing the true disentanglement and editability of our representation. This level of control is a step towards the intuitive, semantic editing capabilities seen in text-driven systems (Wang et al., 2022; Dadashnialehi et al., 2023), bridging the gap between reconstruction and content creation.
Figure 3. Qualitative results on the DTU dataset and demonstration of relighting and editing capabilities. (a) Geometric reconstruction comparison, where our method captures finer details than the baseline. (b) Our reconstructed asset rendered under two novel lighting environments (a warm point light and a cool HDR map), exhibiting physically plausible shading. (c) Material editing demonstration: we modify the albedo map (center) and re-render the object (right), showing the correct appearance update.
4.4. Ablation Studies
To validate the effectiveness of the key components in our MatDecompSDF framework, we conduct a series of ablation studies on the synthetic dataset. We start with our full model and systematically remove or replace individual components, reporting the impact on performance in Table 5.
Table 4. Ablation studies on the synthetic dataset. We report the impact of removing key components from our full MatDecompSDF model on geometry (CD), material decomposition (Albedo PSNR), and perceptual quality (NVS LPIPS).
Variant CD $\downarrow$ Albedo PSNR $\uparrow$ NVS LPIPS $\downarrow$
Ours (Full Model) 0.62 28.45 0.071
w/o Mat. Priors 0.65 24.12 0.085
w/o LPIPS Loss 0.64 27.98 0.112
w/ SH Lighting 0.71 25.03 0.094
w/o PBR Renderer 0.88 - 0.135
Effect of Material Priors
In the first ablation (‘w/o Mat. Priors‘), we remove the material smoothness loss ($\mathcal{L}_{mat\_smooth}$) and the metallic sparsity prior. As the results show, this leads to a significant drop in the PSNR of the decomposed material maps, especially the albedo. Without these priors, the optimization has a strong tendency to bake lighting effects (e.g., soft shadows) into the albedo texture, as it provides an easy way to explain color variations in the input images. This confirms that these physically-motivated priors are crucial for regularizing the ill-posed problem and achieving clean disentanglement.
Effect of Perceptual Loss
Next, we remove the LPIPS term from our photometric loss (‘w/o LPIPS‘), relying solely on the L1 loss. While the PSNR for novel view synthesis remains comparable, the LPIPS score degrades noticeably. Visually, the resulting renderings appear slightly blurrier and lose some of the fine textural detail. This highlights the importance of perceptual metrics in guiding the optimization towards results that are more aligned with human vision, a consideration that is central to creating compelling content for applications from sign language production (Tang et al., 2022) to visual entertainment.
Effect of Lighting Representation
We then replace our flexible MLP-based lighting model with a simpler, 9-coefficient Spherical Harmonics (SH) model (‘w/ SH Lighting‘), similar to that used in early works (Zhou et al., 2021). The performance drops across all metrics. An SH model with low-order coefficients is unable to represent high-frequency lighting, such as sharp shadows or specular highlights from small, bright light sources. The network compensates for this modeling error by baking the residual high-frequency lighting effects into the material maps, leading to inaccurate decomposition. This demonstrates the necessity of a sufficiently expressive lighting representation, a finding consistent with more recent inverse rendering works (Zhang et al., 2022c).
Effect of PBR Renderer
Finally, in the most extreme ablation (‘w/o PBR‘), we replace our entire differentiable PBR shading module with a simple view-dependent appearance model, similar to the original NeRF. In this configuration, the material network $f_{mat}$ effectively becomes a black box that outputs a view-dependent color. As expected, the model completely loses the ability to decompose materials, and the concept of relighting becomes meaningless. While it can still perform novel view synthesis reasonably well, this ablation confirms that a physically-based, differentiable renderer is the indispensable core of our framework and any system aiming for true inverse rendering.
Table 5. Ablation studies on the synthetic dataset. We report the impact of removing key components from our full MatDecompSDF model on geometry (CD) and material decomposition (Albedo PSNR).
4.5. Discussion and Limitations
Our experiments comprehensively demonstrate that MatDecompSDF achieves state-of-the-art performance in jointly reconstructing 3D geometry and decomposing PBR materials from multi-view images. The combination of a high-fidelity implicit surface representation, a physically-based differentiable renderer, and a set of well-designed regularization terms allows our method to achieve a superior level of disentanglement compared to prior work. The ability to produce editable, relightable assets directly from images has significant implications for streamlining content creation pipelines in various industries.
However, our method is not without limitations. First, our current PBR model, while effective for a wide range of common materials, does not explicitly handle more complex optical phenomena. As shown in Figure 4, our framework struggles with highly transparent or translucent objects (e.g., glass, jade) and materials with complex microstructures that produce iridescent or anisotropic effects (e.g., brushed metal, certain fabrics). Modeling these requires more sophisticated light transport simulation and material representations, as explored in specialized works like (Yu et al., 2023) and (Shen et al., 2023), which represent promising avenues for future work.
Figure 4. Limitations and failure cases of our method. (Top) For highly transparent objects like glass, our framework fails to model light transport correctly, resulting in a distorted, opaque reconstruction. (Bottom) For materials with strong anisotropic reflectance like brushed metal, our PBR model cannot capture the direction-dependent highlights, leading to a blurry and inaccurate appearance.
Second, like most neural inverse rendering methods, the performance of MatDecompSDF can degrade when the input views are very sparse or cover only a limited portion of the object. While our priors provide some robustness, significant ambiguity remains in unobserved regions. Integrating generative priors from large-scale models, perhaps through techniques like variational score distillation (Wang et al., 2023), could help to plausibly complete the shape and materials in such data-scarce scenarios.
Finally, as 3D capture technology becomes more democratized, it will increasingly rely on data from distributed, personal devices. This raises important questions about privacy and security. The ability to reconstruct detailed 3D models could be misused, and the underlying sensing systems could be vulnerable, a concern echoed in the domain of WiFi sensing which has demonstrated risks like keystroke sniffing (Huang et al., 2024a). Future work should explore privacy-preserving reconstruction techniques, potentially leveraging federated learning frameworks (Liu et al., 2024a) to train models without centralizing sensitive user data, and ensuring the security of the underlying capture and processing pipeline.

### v2

@@
 1. Introduction
-The demand for high-fidelity, interactive, and editable 3D content has become a cornerstone of modern digital experiences, fueling advancements in virtual and augmented reality, digital twins, and the entertainment industry (Guo et al., 2023).
+The demand for high-fidelity, interactive, and editable 3D content has become a cornerstone of modern digital experiences, fueling advancements in virtual and augmented reality, digital twins, and the entertainment industry .
 Traditionally, creating such photorealistic digital assets is a labor-intensive process, often requiring expert artists to manually model geometry and define complex material properties.
 The advent of neural scene representations has marked a paradigm shift, offering a new avenue for synthesizing and reconstructing the 3D world directly from images.
-Groundbreaking work on Neural Radiance Fields (NeRF) demonstrated an unprecedented ability to synthesize photorealistic novel views from a collection of images by representing a scene as a continuous volumetric function (Mildenhall et al., 2020).
-This seminal work spurred a wave of research exploring dynamic scenes (Pumarola et al., 2021; Li et al., 2021b) and even articulated subjects like humans (Liu et al., 2021; Li et al., 2023; Wu et al., 2023).
+Groundbreaking work on Neural Radiance Fields (NeRF) demonstrated an unprecedented ability to synthesize photorealistic novel views from a collection of images by representing a scene as a continuous volumetric function (mildenhall2020nerf).
+This seminal work spurred a wave of research exploring dynamic scenes (pumarola2021dnerf; li2021realtimedynamic) and even articulated subjects like humans (li2023tava; wu2023magicpony).
 However, the implicit volumetric nature of NeRF makes it challenging to extract explicit, high-quality surface geometry, which is crucial for most graphics pipelines.
 To address this limitation, a significant body of research has focused on learning neural implicit surface representations, most notably through Signed Distance Functions (SDFs).
-Frameworks like NeuS (Wang et al., 2021) and VolSDF (Yariv et al., 2021) ingeniously bridge the gap between volume rendering and surface rendering, enabling the extraction of detailed and coherent meshes from multi-view images.
-The pursuit of efficiency and detail in these reconstructions continues, with methods exploring explicit representations like tensor decompositions (Jin et al., 2022; Fridovich-Keil et al., 2023) and optimization ”baking” strategies to accelerate rendering (Chen et al., 2023a).
-This hierarchical refinement of detail, from coarse volumetric fields to precise surfaces, mirrors the effectiveness of hierarchical models in other domains, such as understanding the nested structure of language in sign language translation (Guo et al., 2018, 2019b) or parsing complex scenes in visual dialog (Guo et al., 2020, 2021).
+Frameworks like NeuS (wang2021neus) and VolSDF (yariv2021volsdf) ingeniously bridge the gap between volume rendering and surface rendering, enabling the extraction of detailed and coherent meshes from multi-view images.
+The pursuit of efficiency and detail in these reconstructions continues, with methods exploring explicit representations like tensor decompositions (jin2022tensoir; fridovich2023kplanes) and optimization ”baking” strategies to accelerate rendering (chen2023bakesdf).
+This hierarchical refinement of detail, from coarse volumetric fields to precise surfaces, mirrors the effectiveness of hierarchical models in other domains, such as understanding the nested structure of language in sign language translation or parsing complex scenes in visual dialog.
 Nevertheless, geometry alone is insufficient for true photorealism.
@@
 The ultimate goal is thus inverse rendering: the challenging, ill-posed problem of disentangling these coupled factors from 2D observations.
-Early explorations in the neural domain sought to factorize NeRF into reflectance and illumination components (Zhou et al., 2021; Boss et al., 2021), often under simplified assumptions.
+Early explorations in the neural domain sought to factorize NeRF into reflectance and illumination components (zhou2021nerfactor; boss2021nerd), often under simplified assumptions.
 Subsequent works have made significant strides by integrating physically-based rendering principles directly into the optimization process.
-These methods aim to decompose appearance into standardized Physically-Based Rendering (PBR) parameters—such as albedo, roughness, and metallic—by leveraging differentiable renderers (Zhang et al., 2021b, 2022b).
-This disentanglement is critical, as failing to separate these components can lead to ambiguity, a challenge analogous to disambiguating events from correlated signals in audio-visual analysis (Zhou et al., 2022) or grounding language in proposal-free video understanding (Li et al., 2021a).
-State-of-the-art inverse rendering frameworks have achieved remarkable results, capable of extracting high-quality triangular meshes along with their PBR material maps and complex lighting environments (Munkberg et al., 2022; Zhang et al., 2023).
-These methods have been extended to handle complex indoor scenes (Yen-Chen et al., 2021; Chang et al., 2023), unstructured ”in-the-wild” online image collections (Liu et al., 2022), and even animatable human avatars with relightable properties (Zhao et al., 2022).
-The development of comprehensive frameworks has further unified these disparate efforts (Huang et al., 2023b).
+These methods aim to decompose appearance into standardized Physically-Based Rendering (PBR) parameters—such as albedo, roughness, and metallic—by leveraging differentiable renderers (zhang2021physg; zhang2022iron).
+This disentanglement is critical, as failing to separate these components can lead to ambiguity, a challenge analogous to disambiguating events from correlated signals in audio-visual analysis or grounding language in proposal-free video understanding .
+State-of-the-art inverse rendering frameworks have achieved remarkable results, capable of extracting high-quality triangular meshes along with their PBR material maps and complex lighting environments (munkberg2022nvdiffrec; zhang2023invrender).
+These methods have been extended to handle complex indoor scenes (yen2021nerfilab; chang2023neilnet), unstructured ”in-the-wild” online image collections (liu2022neroic), and even animatable human avatars with relightable properties (zhao2022relightableavatar).
+The development of comprehensive frameworks has further unified these disparate efforts.
 Despite this progress, significant challenges remain.
-First, many high-fidelity methods are computationally expensive, hindering their practical application; recent advances in real-time relighting with 3D Gaussians (Huang et al., 2024c) and efficient representations (Chen et al., 2022) are beginning to address this.
-The need for efficient models is a universal concern, echoing the importance of model compression techniques like multi-objective convex quantization in deploying large networks (Fan et al., 2024).
-Second, most models are limited to simple, opaque materials, struggling with optically complex phenomena like translucency (Yu et al., 2023), anisotropic reflectance (Shen et al., 2023), or the intricate structure of hair (Chen et al., 2023b).
-Third, robustly handling diverse real-world data requires sophisticated techniques, from modeling visibility and occlusion (Zhang et al., 2021a; Srinivasan et al., 2023) to leveraging specialized information like polarization (Mochida et al., 2022).
+First, many high-fidelity methods are computationally expensive, hindering their practical application; recent advances in real-time relighting with 3D Gaussians (huang2024relightable3dgaussian) and efficient representations (chen2022splitgaussian) are beginning to address this.
+The need for efficient models is a universal concern, echoing the importance of model compression techniques like multi-objective convex quantization in deploying large networks .
+Second, most models are limited to simple, opaque materials, struggling with optically complex phenomena like translucency (yu2023neuraltransmitted), anisotropic reflectance (shen2023neraf), or the intricate structure of hair (chen2023neuralstrands).
+Third, robustly handling diverse real-world data requires sophisticated techniques, from modeling visibility and occlusion (zhang2021nerfv; srinivasan2023nero) to leveraging specialized information like polarization (mochida2022polarizednerf).
 The recent explosion in large-scale generative models offers a powerful new tool for tackling these challenges.
-Text-to-2D diffusion models have been ingeniously repurposed as powerful 3D priors through score distillation sampling (Poole et al., 2023), enabling text-to-3D content creation (Lin et al., 2023).
-These generative techniques are evolving to produce higher-fidelity and more diverse results (Wang et al., 2023) and are increasingly capable of disentangling geometry and appearance in a text-guided manner (Chen et al., 2023c).
-This allows for semantic texturing (Richardson et al., 2023) and material assignment (Seo et al., 2023), where textual concepts like ”metallic” or ”wooden” are translated into PBR values, a process informed by material-aware image generation (Xu et al., 2023).
-Furthermore, multi-modal models like CLIP facilitate direct, text-driven editing of neural scenes (Wang et al., 2022), enabling intuitive control over object materials (Dadashnialehi et al., 2023) and textures (Baatz et al., 2022).
-This convergence of vision, language, and graphics points towards a future of highly controllable and semantic 3D content creation, akin to how cross-modal understanding enhances tasks like sign language production (Tang et al., 2022) and crowd counting (Guo et al., 2019a).
+Text-to-2D diffusion models have been ingeniously repurposed as powerful 3D priors through score distillation sampling (poole2022dreamfusion), enabling text-to-3D content creation (lin2023magic3d).
+These generative techniques are evolving to produce higher-fidelity and more diverse results (wang2023prolificdreamer) and are increasingly capable of disentangling geometry and appearance in a text-guided manner (chen2023fantasia3d).
+This allows for semantic texturing (richardson2023texture) and material assignment (seo2023texfusion), where textual concepts like ”metallic” or ”wooden” are translated into PBR values, a process informed by material-aware image generation (xu2023materialaware).
+Furthermore, multi-modal models like CLIP facilitate direct, text-driven editing of neural scenes (wang2022clipnerf), enabling intuitive control over object materials (dadashnialehi2023neuralscenedecoration) and textures (baatz2022nerftex).
+This convergence of vision, language, and graphics points towards a future of highly controllable and semantic 3D content creation, akin to how cross-modal understanding enhances tasks like sign language production and crowd counting .
 Beyond the visual spectrum, the broader field of sensing is also moving towards extracting nuanced information from ubiquitous signals.
-The ability to recognize human activities (Huang et al., 2020, 2023a), gestures (Gu et al., 2023a), emotions (Gu et al., 2023b), and even vital pulmonary functions (Zhao et al., 2025; Wang et al., 2024) from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements.
-The security and privacy implications of such advanced sensing, from fingerprinting attacks (Huang et al., 2024b) to keystroke eavesdropping (Huang et al., 2024a), highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data.
-Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search (Liu et al., 2024b) and experience-driven model migration (Liu et al., 2024a), provide a roadmap for future privacy-preserving 3D capture ecosystems.
+The ability to recognize human activities (huang2023phaseanti), gestures (gu2023attentionbased), emotions (gu2023wife), and even vital pulmonary functions (zhao2025wipulmo; wang2024targetoriented) from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements.
+The security and privacy implications of such advanced sensing, from fingerprinting attacks (huang2024phyfinatt) to keystroke eavesdropping (huang2024keystrokesniffer), highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data.
+Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search (liu2024finch) and experience-driven model migration (liu2024federated), provide a roadmap for future privacy-preserving 3D capture ecosystems.
 This thesis situates itself within this rich context, aiming to push the boundaries of high-fidelity, physically-based inverse rendering of general objects.
-We focus on robustly and efficiently disentangling geometry and complex material properties from multi-view images, drawing inspiration from both the explicit structure of physically-based rendering and the powerful priors of modern neural networks (Wu et al., 2024).
+We focus on robustly and efficiently disentangling geometry and complex material properties from multi-view images, drawing inspiration from both the explicit structure of physically-based rendering and the powerful priors of modern neural networks (wu2024extracting).
 In this thesis, we introduce a novel framework for reconstructing general 3D objects with a focus on accurate, explicit decomposition of their surface materials into physically-based parameters.
@@
 2.1. Neural Scene Representations for Geometry
-The foundation of modern neural rendering was laid by Neural Radiance Fields (NeRF) (Mildenhall et al., 2020), which represents a scene’s geometry and view-dependent appearance using a multilayer perceptron (MLP).
+The foundation of modern neural rendering was laid by Neural Radiance Fields (NeRF) (mildenhall2020nerf), which represents a scene’s geometry and view-dependent appearance using a multilayer perceptron (MLP).
 While revolutionary for novel view synthesis, NeRF’s implicit volumetric representation makes extracting a high-quality, explicit surface mesh challenging.
 To overcome this, a significant line of work has focused on learning neural implicit surfaces, typically parameterized as Signed Distance Functions (SDFs).
-VolSDF (Yariv et al., 2021) and NeuS (Wang et al., 2021) were pioneering works in this direction, proposing novel formulations to convert SDFs into volume densities for rendering, thereby enabling the optimization of a coherent surface from 2D images.
+VolSDF (yariv2021volsdf) and NeuS (wang2021neus) were pioneering works in this direction, proposing novel formulations to convert SDFs into volume densities for rendering, thereby enabling the optimization of a coherent surface from 2D images.
 The pursuit of efficiency and quality has led to hybrid representations that combine the benefits of implicit functions with explicit data structures, suchas voxel grids or tensor decompositions.
-For instance, TensoIR (Jin et al., 2022) and K-Planes (Fridovich-Keil et al., 2023) utilize explicit low-rank tensor representations to achieve significantly faster training and rendering speeds.
-Other methods focus on ”baking” a trained neural field into an explicit structure like a mesh with features, enabling real-time rendering post-optimization (Chen et al., 2023a).
-The recent introduction of 3D Gaussian Splatting (Chen et al., 2022) has offered a powerful alternative, representing scenes as a collection of anisotropic Gaussians that can be rendered in real-time with high fidelity.
-The challenge of efficiently representing complex data is not unique to 3D vision; similar concerns motivate research in model compression (Fan et al., 2024) and the design of efficient architectures for tasks ranging from visual dialog (Guo et al., 2020) to crowd counting (Guo et al., 2019a).
+For instance, TensoIR (jin2022tensoir) and K-Planes (fridovich2023kplanes) utilize explicit low-rank tensor representations to achieve significantly faster training and rendering speeds.
+Other methods focus on ”baking” a trained neural field into an explicit structure like a mesh with features, enabling real-time rendering post-optimization (chen2023bakesdf).
+The recent introduction of 3D Gaussian Splatting (chen2022splitgaussian) has offered a powerful alternative, representing scenes as a collection of anisotropic Gaussians that can be rendered in real-time with high fidelity.
+The challenge of efficiently representing complex data is not unique to 3D vision; similar concerns motivate research in model compression and the design of efficient architectures for tasks ranging from visual dialog to crowd counting .
 2.2. Neural Inverse Rendering
@@
 Early methods extended the NeRF framework to jointly estimate reflectance and illumination.
-NeRFactor (Zhou et al., 2021) factorized the scene into a shape, a BRDF, and an illumination model represented by spherical harmonics.
-NeRD (Boss et al., 2021) further explored reflectance decomposition from challenging in-the-wild images.
+NeRFactor (zhou2021nerfactor) factorized the scene into a shape, a BRDF, and an illumination model represented by spherical harmonics.
+NeRD (boss2021nerd) further explored reflectance decomposition from challenging in-the-wild images.
 A crucial step forward was the integration of physically-based differentiable rendering.
 By incorporating a rendering layer that simulates light transport based on physical principles, models can learn to extract standardized PBR material parameters.
-PhySG (Zhang et al., 2021b) and IRON (Zhang et al., 2022b) were influential in demonstrating the joint optimization of neural SDFs with material properties like diffuse albedo and roughness.
-The current state-of-the-art is represented by powerful frameworks like NVDIFFREC (Munkberg et al., 2022), which directly optimizes a triangle mesh, its PBR materials, and an HDR light map to match input images.
-Mat-NeuS (Zhang et al., 2022a) extends the high-fidelity surface reconstruction of NeuS to simultaneously acquire material properties.
-Neural-PIL (Boss et al., 2022) introduced pre-integrated lighting to efficiently handle complex reflectance.
-These methods have been successfully applied to reconstruct general objects (Zhang et al., 2023) and even full indoor scenes (Yen-Chen et al., 2021; neil-net:2023).
+PhySG (zhang2021physg) and IRON (zhang2022iron) were influential in demonstrating the joint optimization of neural SDFs with material properties like diffuse albedo and roughness.
+The current state-of-the-art is represented by powerful frameworks like NVDIFFREC (munkberg2022nvdiffrec), which directly optimizes a triangle mesh, its PBR materials, and an HDR light map to match input images.
+Mat-NeuS (zhang2022matneus) extends the high-fidelity surface reconstruction of NeuS to simultaneously acquire material properties.
+Neural-PIL (boss2022neuralpil) introduced pre-integrated lighting to efficiently handle complex reflectance.
+These methods have been successfully applied to reconstruct general objects (zhang2023invrender) and even full indoor scenes (yen2021nerfilab).
 The core challenge in all these methods is robust disentanglement, which requires careful regularization and priors.
-This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition (Huang et al., 2020, 2023a) or building secure authentication systems resilient to fingerprinting attacks (Huang et al., 2024b).
+This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition (huang2023phaseanti) or building secure authentication systems resilient to fingerprinting attacks (huang2024phyfinatt).
 Our work contributes to this line of research by proposing new techniques to improve the robustness of material decomposition under unknown lighting.
-We also draw inspiration from how hierarchical models can parse complex structures, a principle proven effective in tasks like sign language translation (Guo et al., 2018, 2019b).
+We also draw inspiration from how hierarchical models can parse complex structures, a principle proven effective in tasks like sign language translation.
 2.3. Generative Priors for 3D Content Creation
 The rise of large-scale generative models, particularly 2D diffusion models, has unlocked new paradigms for 3D content creation.
-DreamFusion (Poole et al., 2023) introduced Score Distillation Sampling (SDS), a method to ”distill” the knowledge of a pre-trained 2D text-to-image model to guide the optimization of a 3D representation like NeRF.
-This spawned a flurry of research in text-to-3D generation, with follow-up works like Magic3D (Lin et al., 2023) improving resolution and efficiency.
-ProlificDreamer (Wang et al., 2023) proposed Variational Score Distillation (VSD) to enhance the diversity and fidelity of generated results.
+DreamFusion (poole2022dreamfusion) introduced Score Distillation Sampling (SDS), a method to ”distill” the knowledge of a pre-trained 2D text-to-image model to guide the optimization of a 3D representation like NeRF.
+This spawned a flurry of research in text-to-3D generation, with follow-up works like Magic3D (lin2023magic3d) improving resolution and efficiency.
+ProlificDreamer (wang2023prolificdreamer) proposed Variational Score Distillation (VSD) to enhance the diversity and fidelity of generated results.
 A key focus of recent work is the disentanglement of geometry and appearance within this generative framework.
-Fantasia3D (Chen et al., 2023c) and TexFusion (Seo et al., 2023) explicitly aim to generate not just a shape, but also assign meaningful textures and PBR materials based on textual prompts.
+Fantasia3D (chen2023fantasia3d) and TexFusion (seo2023texfusion) explicitly aim to generate not just a shape, but also assign meaningful textures and PBR materials based on textual prompts.
 This leverages the powerful multi-modal understanding of joint vision-language models like CLIP.
-Such models are also used for text-guided editing of existing 3D shapes (Richardson et al., 2023) or neural fields (Wang et al., 2022), enabling intuitive manipulation of scene properties (Dadashnialehi et al., 2023).
-The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog (Guo et al., 2021), video grounding (Li et al., 2021a), and even cross-modal emotion recognition (Gu et al., 2023b), all of which seek to bridge the gap between different data modalities.
+Such models are also used for text-guided editing of existing 3D shapes (richardson2023texture) or neural fields (wang2022clipnerf), enabling intuitive manipulation of scene properties (dadashnialehi2023neuralscenedecoration).
+The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog , video grounding , and even cross-modal emotion recognition (gu2023wife), all of which seek to bridge the gap between different data modalities.
 2.4. Modeling of Advanced Materials and Dynamic Scenes
@@
 For complex materials, specialized representations are needed.
-NeRF-V (Zhang et al., 2021a) and NeRO (Srinivasan et al., 2023) explicitly model visibility and occlusion to handle complex shadowing and relighting.
-Polarized-NeRF (Mochida et al., 2022) leverages polarization cues to better separate diffuse and specular components.
-For translucent objects with subsurface scattering, methods like Neural Transmitted Radiance Fields (Yu et al., 2023) model the path of light through the object’s interior.
-Other works tackle anisotropic materials (Shen et al., 2023) or highly complex microstructures like hair (Chen et al., 2023b).
-The challenge of representing complex appearance is also addressed by works that learn decoupled neural textures (Baatz et al., 2022) or extract neural BRDFs (Wu et al., 2024).
+NeRF-V (zhang2021nerfv) and NeRO (srinivasan2023nero) explicitly model visibility and occlusion to handle complex shadowing and relighting.
+Polarized-NeRF (mochida2022polarizednerf) leverages polarization cues to better separate diffuse and specular components.
+For translucent objects with subsurface scattering, methods like Neural Transmitted Radiance Fields (yu2023neuraltransmitted) model the path of light through the object’s interior.
+Other works tackle anisotropic materials (shen2023neraf) or highly complex microstructures like hair (chen2023neuralstrands).
+The challenge of representing complex appearance is also addressed by works that learn decoupled neural textures (baatz2022nerftex) or extract neural BRDFs (wu2024extracting).
 Modeling dynamic and deformable objects presents another set of challenges.
-D-NeRF (Pumarola et al., 2021) pioneered the use of a deformation field to map a dynamic scene to a canonical space.
-This concept has been extended to create controllable and animatable avatars of human actors (Liu et al., 2021; Li et al., 2023), often incorporating relighting capabilities (Zhao et al., 2022).
-The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi (Gu et al., 2023a).
-The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (Liu et al., 2022) or from distributed agents in a federated learning setting (Liu et al., 2024b, a).
-The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems (Huang et al., 2024a).
-Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis (Zhou et al., 2022) or respiratory healthcare sensing (Zhao et al., 2025; Wang et al., 2024), is key.
+D-NeRF (pumarola2021dnerf) pioneered the use of a deformation field to map a dynamic scene to a canonical space.
+This concept has been extended to create controllable and animatable avatars of human actors , often incorporating relighting capabilities (zhao2022relightableavatar).
+The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi (gu2023attentionbased).
+The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (liu2022neroic) or from distributed agents in a federated learning setting (liu2024finch; liu2024federated).
+The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems (huang2024keystrokesniffer).
+Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis or respiratory healthcare sensing (zhao2025wipulmo; wang2024targetoriented), is key.
 3. Methodology
@@
 The discrepancy between the rendered colors and the input images drives the joint optimization of all three network components.
-This end-to-end optimization is analogous to complex inference processes seen in other domains, such as iterative context-aware graph inference for visual dialog (Guo et al., 2020).
+This end-to-end optimization is analogous to complex inference processes seen in other domains, such as iterative context-aware graph inference for visual dialog .
 Figure 1.
@@
 3.2. Neural Scene Representation
-We represent the scene using three distinct MLPs to effectively disentangle geometry, material, and lighting, a strategy proven more robust than using a single monolithic network (Zhang et al., 2021b).
+We represent the scene using three distinct MLPs to effectively disentangle geometry, material, and lighting, a strategy proven more robust than using a single monolithic network (zhang2021physg).
 3.2.1. Geometry as a Neural SDF
-We represent the object’s geometry using a neural network $f_{geo}$ that approximates a Signed Distance Function (SDF).
-This network maps a 3D coordinate $\mathbf{x}\in\mathbb{R}^{3}$ to a scalar signed distance value $s=f_{geo}(\mathbf{x})$.
-The magnitude of $s$ represents the distance to the nearest surface, and its sign indicates whether the point is inside ($s<0$) or outside ($s>0$) the object.
-The object surface $\mathcal{S}$ is thus implicitly defined as the zero-level set of the SDF: $\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{3}\mid f_{geo}(\mathbf{x})=0\}$.
-This representation, popularized by works like NeuS (Wang et al., 2021), allows for the extraction of high-fidelity, watertight meshes, which is a significant advantage over direct volumetric representations like NeRF (Mildenhall et al., 2020).
-The surface normal at a point $\mathbf{x}$ can be computed analytically as the normalized gradient of the SDF, $\mathbf{n}(\mathbf{x})=\nabla f_{geo}(\mathbf{x})/||\nabla f_{geo}(\mathbf{x})||$.
+[…]
 3.2.2. Spatially-Varying PBR Materials
@@
 By conditioning the material properties on the 3D location, we can capture spatially-varying materials, such as objects with different painted parts or worn-out surfaces.
-This explicit decomposition is crucial for downstream applications like material editing, a feature lacking in methods that only model view-dependent appearance (Poole et al., 2023; Lin et al., 2023).
+This explicit decomposition is crucial for downstream applications like material editing, a feature lacking in methods that only model view-dependent appearance (poole2022dreamfusion; lin2023magic3d).
 3.2.3. Environment Lighting
@@
 The lighting is represented by a third MLP, $f_{light}$, which maps an incoming light direction vector $\boldsymbol{\omega}_{i}\in\mathbb{S}^{2}$ to an RGB color value $\mathbf{L}_{i}(\boldsymbol{\omega}_{i})=f_{light}(\boldsymbol{\omega}_{i})$.
-This formulation is flexible enough to represent complex, high-frequency lighting environments, which is a notable improvement over simpler models like spherical harmonics used in some earlier works (Zhou et al., 2021).
+This formulation is flexible enough to represent complex, high-frequency lighting environments, which is a notable improvement over simpler models like spherical harmonics used in some earlier works (zhou2021nerfactor).
 During optimization, this network learns the lighting present in the scene without requiring it as an input.
@@
 The core of our method is a differentiable rendering process that connects our 3D scene representation to the 2D input images.
-We adopt the volume rendering principle from NeuS (Wang et al., 2021) to robustly find the surface intersection and then apply a physically-based rendering equation at that surface point.
+We adopt the volume rendering principle from NeuS (wang2021neus) to robustly find the surface intersection and then apply a physically-based rendering equation at that surface point.
 3.3.1. Ray Marching and Surface Finding
@@
 The color of the ray $\mathbf{C}(\mathbf{r})$ is computed by integrating the color and density along the ray:
-[…]
-[…]
+(1) $\mathbf{C}(\mathbf{r})=\int_{t_{n}}^{t_{f}}T(t)w(t)\mathbf{c}(\mathbf{r}(t),\mathbf{d})dt,$
+where $t_{n},t_{f}$ are near/far bounds, $T(t)=\exp(-\int_{t_{n}}^{t}\sigma(s)ds)$ is the transmittance, $w(t)$ is a weighting function, and $\mathbf{c}(\cdot)$ is the color at point $\mathbf{r}(t)$.
 Instead of a generic density $\sigma$, we use a weighting function $w(t)$ that is guaranteed to be a single, unbiased peak at the surface location.
-Following NeuS (Wang et al., 2021), we define a density-like function $\sigma(t)$ from the SDF value and then derive the weight $w(t)=T(t)\sigma(t)$.
+Following NeuS (wang2021neus), we define a density-like function $\sigma(t)$ from the SDF value and then derive the weight $w(t)=T(t)\sigma(t)$.
 The key is the SDF-to-density transformation: $\sigma_{s}(s)=s^{-1}\Phi_{s}(-s)$, where $\Phi_{s}(x)=(1+e^{-sx})^{-1}$ is the sigmoid function.
@@
 The outgoing radiance $\mathbf{L}_{o}$ from point $\mathbf{x}_{s}$ in the view direction $\boldsymbol{\omega}_{o}=-\mathbf{d}$ is given by an integral over the hemisphere $\Omega$:
-[…]
+(2) $\mathbf{L}_{o}(\mathbf{x}_{s},\boldsymbol{\omega}_{o})=\int_{\Omega}f_{r}(\mathbf{x}_{s},\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o})\mathbf{L}_{i}(\boldsymbol{\omega}_{i})(\mathbf{n}(\mathbf{x}_{s})\cdot\boldsymbol{\omega}_{i})d\boldsymbol{\omega}_{i},$
 where $f_{r}$ is the Bidirectional Reflectance Distribution Function (BRDF), $\mathbf{L}_{i}$ is the incoming radiance from direction $\boldsymbol{\omega}_{i}$, and $\mathbf{n}$ is the surface normal.
@@
 $k_{d},k_{s}$ are diffuse/specular coefficients.
-The terms $D,F,G$ are standard functions of the surface roughness $r$, view/light directions, and the Fresnel term $F$ additionally depends on the metallic property $m$ and albedo $\boldsymbol{\alpha}$ (Munkberg et al., 2022).
+The terms $D,F,G$ are standard functions of the surface roughness $r$, view/light directions, and the Fresnel term $F$ additionally depends on the metallic property $m$ and albedo $\boldsymbol{\alpha}$ (munkberg2022nvdiffrec).
 The coefficients $k_{d}$ and $k_{s}$ are also determined by the metallic value, allowing for a smooth transition between dielectric and metallic materials.
@@
 The networks are trained end-to-end by minimizing a composite loss function.
-The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production (Tang et al., 2022) and WiFi-based sensing (Wang et al., 2024).
+The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production and WiFi-based sensing (wang2024targetoriented).
 3.4.1. Photometric Reconstruction Loss
 The primary supervision signal is the photometric reconstruction loss, which measures the difference between the rendered pixel colors $\mathbf{C}(\mathbf{r})$ and the ground-truth colors from the input images $\mathbf{C}_{gt}(\mathbf{r})$.
-We use a combination of an L1 loss and the LPIPS perceptual loss (Srinivasan et al., 2023):
-(4) $\mathcal{L}_{photo}=\lambda_{L1}||\mathbf{C}(\mathbf{r})-\mathbf{C}_{gt}(\mathbf{r})||_{1}+\lambda_{LPIPS}\mathcal{L}_{LPIPS}(\mathbf{C}(\mathbf{r}),\mathbf{C}_{gt}(\mathbf{r})).$
+We use a combination of an L1 loss and the LPIPS perceptual loss (srinivasan2023nero):
+[…]
 where $\lambda_{L1},\lambda_{LPIPS}$ are loss weights. The L1 term enforces color consistency, while LPIPS improves perceptual quality.
@@
 To ensure that $f_{geo}$ learns a valid SDF, we apply the Eikonal regularization, which encourages the norm of the SDF gradient to be unity everywhere:
-(5) $\mathcal{L}_{eikonal}=\mathbb{E}_{\mathbf{x}}\left(||\nabla f_{geo}(\mathbf{x})||_{2}-1\right)^{2}.$
+[…]
 where the expectation $\mathbb{E}_{\mathbf{x}}$ is taken over points $\mathbf{x}$ sampled in the 3D space.
 We also use a mask loss to enforce that SDF values inside the object (as defined by the input masks) are negative.
-This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains (Huang et al., 2024b).
+This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains (huang2024phyfinatt).
 3.4.3. Material and Lighting Priors
@@
 A material smoothness prior encourages the predicted material properties to be locally smooth on the object’s surface:
-(6) $\mathcal{L}_{mat\_smooth}=\mathbb{E}_{\mathbf{x}_{s}}\left(||\nabla\boldsymbol{\alpha}(\mathbf{x}_{s})||_{1}+||\nabla r(\mathbf{x}_{s})||_{1}\right).$
+[…]
 where $\mathbf{x}_{s}$ are points sampled on the surface, and the gradient is computed over the surface.
@@
 We also regularize the total intensity and smoothness of the predicted environment light map from $f_{light}$ to prevent overly complex lighting from being ”baked” into the material albedo.
-The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning (Liu et al., 2024b) or visual understanding (Guo et al., 2019b).
+The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning (liu2024finch) or visual understanding .
 3.4.4. Total Loss
@@
 3.5. Implementation Details
-Our geometry and material networks are implemented as MLPs with 8 layers, 256 hidden units, and sinusoidal positional encodings (Mildenhall et al., 2020) for the input coordinates to capture high-frequency details.
+Our geometry and material networks are implemented as MLPs with 8 layers, 256 hidden units, and sinusoidal positional encodings (mildenhall2020nerf) for the input coordinates to capture high-frequency details.
 The lighting network is a smaller 4-layer MLP.
@@
 We evaluate our method on both synthetic and real-world datasets to demonstrate its effectiveness and generalization capabilities.
-The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language (Guo et al., 2019b) to ensuring robust performance in WiFi-based activity recognition under interference (Huang et al., 2023a).
+The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language to ensuring robust performance in WiFi-based activity recognition under interference (huang2023phaseanti).
 Synthetic Dataset
-For quantitative evaluation of both geometry and material decomposition, we use a synthetic dataset inspired by the one used in NVDIFFREC (Munkberg et al., 2022).
+For quantitative evaluation of both geometry and material decomposition, we use a synthetic dataset inspired by the one used in NVDIFFREC (munkberg2022nvdiffrec).
 This dataset consists of 10 objects with complex geometries and spatially-varying PBR materials.
@@
 Real-world Dataset
-To validate our method on real-world data, we use the widely-adopted DTU dataset (jensen2014large).
+To validate our method on real-world data, we use the widely-adopted DTU dataset .
 The DTU dataset contains multi-view images of various objects captured under controlled laboratory conditions with a fixed camera rig and several different lighting conditions.
-This dataset is a standard benchmark for multi-view surface reconstruction (Wang et al., 2021; Yariv et al., 2021).
+This dataset is a standard benchmark for multi-view surface reconstruction (wang2021neus; yariv2021volsdf).
 Since ground-truth material properties are not available for these real objects, evaluation is primarily focused on the quality of the geometric reconstruction and the visual fidelity of novel view synthesis and relighting results.
-The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals (Gu et al., 2023a).
+The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals (gu2023attentionbased).
 4.1.2. Evaluation Metrics
 We adopt a comprehensive set of metrics to evaluate different aspects of our framework’s performance.
-The choice of metrics is critical for a fair comparison, reflecting the multi-faceted nature of the inverse rendering problem, much like how multi-objective optimization is key for efficient model compression (Fan et al., 2024).
+The choice of metrics is critical for a fair comparison, reflecting the multi-faceted nature of the inverse rendering problem, much like how multi-objective optimization is key for efficient model compression .
 Geometry Reconstruction Quality
@@
 We evaluate the quality of the appearance and the decomposed materials using several metrics:
-  • Novel View Synthesis: For both datasets, we evaluate the quality of rendered novel views against ground-truth images using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and the Learned Perceptual Image Patch Similarity (LPIPS) (Srinivasan et al., 2023).
-PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog (Guo et al., 2021) or emotion recognition (Gu et al., 2023b).
+  • Novel View Synthesis: For both datasets, we evaluate the quality of rendered novel views against ground-truth images using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and the Learned Perceptual Image Patch Similarity (LPIPS) (srinivasan2023nero).
+PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog or emotion recognition (gu2023wife).
   • Material Map Accuracy: On the synthetic dataset, where ground-truth PBR maps are available, we compute the PSNR between our decomposed material maps (albedo, roughness, metallic) and the ground-truth maps.
@@
 Our primary baselines include:
-  • NeuS (Wang et al., 2021): A state-of-the-art method for high-quality neural surface reconstruction. This baseline represents methods that focus solely on geometry.
-  • NVDIFFREC (Munkberg et al., 2022): A leading method for inverse rendering that directly optimizes a mesh representation. It is a strong baseline for joint geometry and material optimization.
-  • Mat-NeuS (Zhang et al., 2022a): An extension of NeuS that incorporates material decomposition, making it a very relevant baseline for our approach.
-  • IRON (Zhang et al., 2022b): Another powerful inverse rendering framework that demonstrates high-quality object reconstruction and material estimation.
+  • NeuS (wang2021neus): A state-of-the-art method for high-quality neural surface reconstruction. This baseline represents methods that focus solely on geometry.
+  • NVDIFFREC (munkberg2022nvdiffrec): A leading method for inverse rendering that directly optimizes a mesh representation. It is a strong baseline for joint geometry and material optimization.
+  • Mat-NeuS (zhang2022matneus): An extension of NeuS that incorporates material decomposition, making it a very relevant baseline for our approach.
+  • IRON (zhang2022iron): Another powerful inverse rendering framework that demonstrates high-quality object reconstruction and material estimation.
 For all baseline methods, we use the official publicly available code and follow their recommended hyperparameter settings to ensure a fair comparison.
@@
 The loss weights are set empirically, with $\lambda_{L1}=1.0$, $\lambda_{LPIPS}=0.2$, $\lambda_{eikonal}=0.1$, and $\lambda_{mat}=0.01$.
-The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning (Liu et al., 2024b, a) or dealing with the security implications of advanced sensing technologies (Huang et al., 2024a).
+The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning (liu2024finch; liu2024federated) or dealing with the security implications of advanced sensing technologies (huang2024keystrokesniffer).
 Our code and pre-trained models will be made publicly available to facilitate future research.
@@
 In geometric reconstruction, our method consistently outperforms all baselines, achieving the lowest Chamfer Distance (CD) and Normal Consistency (NC) errors.
-Compared to NeuS (Wang et al., 2021), which focuses solely on geometry, our joint optimization of shape and material yields a more accurate surface.
+Compared to NeuS (wang2021neus), which focuses solely on geometry, our joint optimization of shape and material yields a more accurate surface.
 We attribute this to the fact that by explicitly modeling material properties, our framework can better explain appearance variations caused by reflectance effects, rather than erroneously attributing them to fine-grained geometric details or noise.
 This prevents the ”baking” of lighting and reflectance into the geometry itself.
-Our geometric fidelity even surpasses that of leading inverse rendering methods like NVDIFFREC (Munkberg et al., 2022) and Mat-NeuS (Zhang et al., 2022a) which we believe is due to our robust lighting model and the comprehensive set of physical and geometric priors that regularize the optimization process.
-The challenge of disentangling correlated factors is a common theme in machine perception, where, for instance, separating an event signal from noise is key in audio-visual analysis (Zhou et al., 2022).
+Our geometric fidelity even surpasses that of leading inverse rendering methods like NVDIFFREC (munkberg2022nvdiffrec) and Mat-NeuS (zhang2022matneus) which we believe is due to our robust lighting model and the comprehensive set of physical and geometric priors that regularize the optimization process.
+The challenge of disentangling correlated factors is a common theme in machine perception, where, for instance, separating an event signal from noise is key in audio-visual analysis .
 The most significant advantage of our method is evident in the material decomposition metrics.
@@
 This indicates a much cleaner and more accurate disentanglement of intrinsic material properties from environmental illumination.
-While methods like IRON (Zhang et al., 2022b) can recover plausible materials, they often exhibit minor artifacts, such as residual lighting baked into the albedo map.
+While methods like IRON (zhang2022iron) can recover plausible materials, they often exhibit minor artifacts, such as residual lighting baked into the albedo map.
 Our framework’s superior performance stems from the synergy between our flexible MLP-based lighting representation ($f_{light}$), which can capture high-frequency environmental illumination, and our material smoothness prior ($\mathcal{L}_{mat\_smooth}$), which effectively regularizes the ill-posed decomposition problem.
-This demonstrates that a well-constrained optimization, guided by physical principles, is paramount, a lesson that holds true in fields like model compression where multi-objective approaches are needed to balance competing goals (Fan et al., 2024).
+This demonstrates that a well-constrained optimization, guided by physical principles, is paramount, a lesson that holds true in fields like model compression where multi-objective approaches are needed to balance competing goals .
 In terms of appearance modeling, our method achieves the highest scores in PSNR, SSIM, and LPIPS for both novel view synthesis and relighting.
@@
 Quantitative comparison on the synthetic dataset.
-We report Chamfer Distance (CD, mm, $\downarrow$), Normal Consistency (NC, degrees, $\downarrow$), and PSNR (dB, $\uparrow$) for material maps and rendering quality.
+We report Chamfer Distance (CD, $m$, $\downarrow$), Normal Consistency (NC, degrees, $\downarrow$), and PSNR (dB, $\uparrow$) for material maps and rendering quality.
 Best results are in bold.
@@
 CD $\downarrow$ NC $\downarrow$ Albedo Roughness Metallic NVS Relighting
-NeuS (Wang et al., 2021) 0.85 16.2 - - - 29.81 -
-NVDIFFREC (Munkberg et al., 2022) 0.79 15.8 25.12 23.45 24.88 31.23 28.95
-Mat-NeuS (Zhang et al., 2022a) 0.75 15.1 26.03 24.11 25.01 31.55 29.24
-IRON (Zhang et al., 2022b) 0.81 15.5 25.88 23.97 24.92 31.40 29.08
+NeuS (wang2021neus) 0.85 16.2 - - - 29.81 -
+NVDIFFREC (munkberg2022nvdiffrec) 0.79 15.8 25.12 23.45 24.88 31.23 28.95
+Mat-NeuS (zhang2022matneus) 0.75 15.1 26.03 24.11 25.01 31.55 29.24
+IRON (zhang2022iron) 0.81 15.5 25.88 23.97 24.92 31.40 29.08
 Ours (MatDecompSDF) 0.62 13.7 28.45 26.83 27.54 32.89 31.16
@@
 This indicates that our rendered images are not only pixel-wise accurate but are also perceptually closer to the ground-truth views.
-We attribute this to two factors: first, the inclusion of the $\mathcal{L}_{LPIPS}$ term in our photometric loss guides the optimization towards perceptually plausible solutions, a strategy whose importance is recognized in generative modeling (Srinivasan et al., 2023).
+We attribute this to two factors: first, the inclusion of the $\mathcal{L}_{LPIPS}$ term in our photometric loss guides the optimization towards perceptually plausible solutions, a strategy whose importance is recognized in generative modeling (srinivasan2023nero).
 Second, by modeling the scene with physically accurate materials and lighting, our framework produces more realistic shading and highlights, which are key components of perceptual quality.
 This is in contrast to geometry-only methods like NeuS, whose view-dependent appearance model can sometimes produce results that look flat or lack realism under complex lighting.
-The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks (Liu et al., 2024a) to secure WiFi-based gesture recognition (Gu et al., 2023a).
+The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks (liu2024federated) to secure WiFi-based gesture recognition (gu2023attentionbased).
 Table 2. Quantitative comparison for novel view synthesis on the DTU dataset. We report PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$). Best results are in bold.
 Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-NeuS (Wang et al., 2021) 25.31 0.871 0.145
-NVDIFFREC (Munkberg et al., 2022) 25.88 0.885 0.131
-Mat-NeuS (Zhang et al., 2022a) 26.05 0.892 0.124
-IRON (Zhang et al., 2022b) 25.97 0.888 0.129
+NeuS (wang2021neus) 25.31 0.871 0.145
+NVDIFFREC (munkberg2022nvdiffrec) 25.88 0.885 0.131
+Mat-NeuS (zhang2022matneus) 26.05 0.892 0.124
+IRON (zhang2022iron) 25.97 0.888 0.129
 Ours (MatDecompSDF) 26.73 0.906 0.108
-Table 3. Quantitative comparison for novel view synthesis on the DTU dataset. We report PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$). Best results are in bold.
 4.3. Qualitative Comparison
@@
 The visual difference is striking.
-The albedo map produced by our MatDecompSDF is significantly cleaner, showing pure surface colors without the shadowy artifacts present in the results of NVDIFFREC (Munkberg et al., 2022) and Mat-NeuS (Zhang et al., 2022a).
+The albedo map produced by our MatDecompSDF is significantly cleaner, showing pure surface colors without the shadowy artifacts present in the results of NVDIFFREC (munkberg2022nvdiffrec) and Mat-NeuS (zhang2022matneus).
 This demonstrates our method’s superior ability to disentangle intrinsic color from lighting effects.
@@
 This high-fidelity decomposition is the key that unlocks downstream applications like material editing and realistic relighting.
-The clarity of our decomposition can be likened to the goal of achieving interference-independent feature extraction in WiFi sensing (Huang et al., 2023a).
+The clarity of our decomposition can be likened to the goal of achieving interference-independent feature extraction in WiFi sensing (huang2023phaseanti).
 Figure 2.
 Qualitative comparison of PBR material decomposition on the synthetic dataset.
-From top to bottom: Ground Truth, our MatDecompSDF, NVDIFFREC (Munkberg et al., 2022), and Mat-NeuS (Zhang et al., 2022a).
+From top to bottom: Ground Truth, our MatDecompSDF, NVDIFFREC (munkberg2022nvdiffrec), and Mat-NeuS (zhang2022matneus).
 From left to right: Albedo, Roughness, and Metallic maps.
@@
 Geometric Reconstruction
-As shown in Figure LABEL:fig:qualitative_dtu, our method reconstructs 3D shapes with higher geometric fidelity and fewer artifacts compared to the baselines on the DTU dataset.
+Our method r econstructs 3D shapes with higher geometric fidelity and fewer artifacts compared to the baselines on the DTU dataset.
 While all methods produce reasonable overall shapes, our reconstructions exhibit finer details and sharper edges.
@@
 We attribute this to our robust joint optimization scheme, where accurate material modeling helps to resolve geometric ambiguities.
-This principle of leveraging one modality or aspect to disambiguate another is a powerful concept, seen also in works that combine visual and textual information for dialog (Guo et al., 2021) or audio and visual cues for event localization (Zhou et al., 2022).
+This principle of leveraging one modality or aspect to disambiguate another is a powerful concept, seen also in works that combine visual and textual information for dialog or audio and visual cues for event localization .
 Relighting and Material Editing
@@
 The resulting rendering correctly reflects this change, showcasing the true disentanglement and editability of our representation.
-This level of control is a step towards the intuitive, semantic editing capabilities seen in text-driven systems (Wang et al., 2022; Dadashnialehi et al., 2023), bridging the gap between reconstruction and content creation.
+This level of control is a step towards the intuitive, semantic editing capabilities seen in text-driven systems (wang2022clipnerf; dadashnialehi2023neuralscenedecoration), bridging the gap between reconstruction and content creation.
 Figure 3.
@@
 Visually, the resulting renderings appear slightly blurrier and lose some of the fine textural detail.
-This highlights the importance of perceptual metrics in guiding the optimization towards results that are more aligned with human vision, a consideration that is central to creating compelling content for applications from sign language production (Tang et al., 2022) to visual entertainment.
+This highlights the importance of perceptual metrics in guiding the optimization towards results that are more aligned with human vision, a consideration that is central to creating compelling content for applications from sign language production to visual entertainment.
 Effect of Lighting Representation
-We then replace our flexible MLP-based lighting model with a simpler, 9-coefficient Spherical Harmonics (SH) model (‘w/ SH Lighting‘), similar to that used in early works (Zhou et al., 2021).
+We then replace our flexible MLP-based lighting model with a simpler, 9-coefficient Spherical Harmonics (SH) model (‘w/ SH Lighting‘), similar to that used in early works (zhou2021nerfactor).
 The performance drops across all metrics.
@@
 The network compensates for this modeling error by baking the residual high-frequency lighting effects into the material maps, leading to inaccurate decomposition.
-This demonstrates the necessity of a sufficiently expressive lighting representation, a finding consistent with more recent inverse rendering works (Zhang et al., 2022c).
+This demonstrates the necessity of a sufficiently expressive lighting representation, a finding consistent with more recent inverse rendering works (zhang2022neilf).
 Effect of PBR Renderer
@@
 While it can still perform novel view synthesis reasonably well, this ablation confirms that a physically-based, differentiable renderer is the indispensable core of our framework and any system aiming for true inverse rendering.
-Table 5. Ablation studies on the synthetic dataset. We report the impact of removing key components from our full MatDecompSDF model on geometry (CD) and material decomposition (Albedo PSNR).
 4.5. Discussion and Limitations
@@
 As shown in Figure 4, our framework struggles with highly transparent or translucent objects (e.g., glass, jade) and materials with complex microstructures that produce iridescent or anisotropic effects (e.g., brushed metal, certain fabrics).
-Modeling these requires more sophisticated light transport simulation and material representations, as explored in specialized works like (Yu et al., 2023) and (Shen et al., 2023), which represent promising avenues for future work.
+Modeling these requires more sophisticated light transport simulation and material representations, as explored in specialized works like (yu2023neuraltransmitted) and (shen2023neraf), which represent promising avenues for future work.
 Figure 4.
@@
 While our priors provide some robustness, significant ambiguity remains in unobserved regions.
-Integrating generative priors from large-scale models, perhaps through techniques like variational score distillation (Wang et al., 2023), could help to plausibly complete the shape and materials in such data-scarce scenarios.
+Integrating generative priors from large-scale models, perhaps through techniques like variational score distillation (wang2023prolificdreamer), could help to plausibly complete the shape and materials in such data-scarce scenarios.
 Finally, as 3D capture technology becomes more democratized, it will increasingly rely on data from distributed, personal devices.
 This raises important questions about privacy and security.
-The ability to reconstruct detailed 3D models could be misused, and the underlying sensing systems could be vulnerable, a concern echoed in the domain of WiFi sensing which has demonstrated risks like keystroke sniffing (Huang et al., 2024a).
-Future work should explore privacy-preserving reconstruction techniques, potentially leveraging federated learning frameworks (Liu et al., 2024a) to train models without centralizing sensitive user data, and ensuring the security of the underlying capture and processing pipeline.
+The ability to reconstruct detailed 3D models could be misused, and the underlying sensing systems could be vulnerable, a concern echoed in the domain of WiFi sensing which has demonstrated risks like keystroke sniffing (huang2024keystrokesniffer).
+Future work should explore privacy-preserving reconstruction techniques, potentially leveraging federated learning frameworks (liu2024federated) to train models without centralizing sensitive user data, and ensuring the security of the underlying capture and processing pipeline.

### v3

@@
 Groundbreaking work on Neural Radiance Fields (NeRF) demonstrated an unprecedented ability to synthesize photorealistic novel views from a collection of images by representing a scene as a continuous volumetric function (mildenhall2020nerf).
-This seminal work spurred a wave of research exploring dynamic scenes (pumarola2021dnerf; li2021realtimedynamic) and even articulated subjects like humans (li2023tava; wu2023magicpony).
+This seminal work spurred a wave of research exploring dynamic scenes (pumarola2021dnerf) and even articulated subjects like humans (li2023tava; wu2023magicpony).
 However, the implicit volumetric nature of NeRF makes it challenging to extract explicit, high-quality surface geometry, which is crucial for most graphics pipelines.
@@
 Beyond the visual spectrum, the broader field of sensing is also moving towards extracting nuanced information from ubiquitous signals.
-The ability to recognize human activities (huang2023phaseanti), gestures (gu2023attentionbased), emotions (gu2023wife), and even vital pulmonary functions (zhao2025wipulmo; wang2024targetoriented) from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements.
-The security and privacy implications of such advanced sensing, from fingerprinting attacks (huang2024phyfinatt) to keystroke eavesdropping (huang2024keystrokesniffer), highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data.
-Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search (liu2024finch) and experience-driven model migration (liu2024federated), provide a roadmap for future privacy-preserving 3D capture ecosystems.
+The ability to recognize human activities , gestures , emotions , and even vital pulmonary functions from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements.
+The security and privacy implications of such advanced sensing, from fingerprinting attacks to keystroke eavesdropping , highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data.
+Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search and experience-driven model migration , provide a roadmap for future privacy-preserving 3D capture ecosystems.
 This thesis situates itself within this rich context, aiming to push the boundaries of high-fidelity, physically-based inverse rendering of general objects.
@@
 The core challenge in all these methods is robust disentanglement, which requires careful regularization and priors.
-This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition (huang2023phaseanti) or building secure authentication systems resilient to fingerprinting attacks (huang2024phyfinatt).
+This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition or building secure authentication systems resilient to fingerprinting attacks .
 Our work contributes to this line of research by proposing new techniques to improve the robustness of material decomposition under unknown lighting.
@@
 Such models are also used for text-guided editing of existing 3D shapes (richardson2023texture) or neural fields (wang2022clipnerf), enabling intuitive manipulation of scene properties (dadashnialehi2023neuralscenedecoration).
-The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog , video grounding , and even cross-modal emotion recognition (gu2023wife), all of which seek to bridge the gap between different data modalities.
+The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog , video grounding , and even cross-modal emotion recognition , all of which seek to bridge the gap between different data modalities.
 2.4. Modeling of Advanced Materials and Dynamic Scenes
@@
 This concept has been extended to create controllable and animatable avatars of human actors , often incorporating relighting capabilities (zhao2022relightableavatar).
-The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi (gu2023attentionbased).
-The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (liu2022neroic) or from distributed agents in a federated learning setting (liu2024finch; liu2024federated).
-The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems (huang2024keystrokesniffer).
-Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis or respiratory healthcare sensing (zhao2025wipulmo; wang2024targetoriented), is key.
+The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi .
+The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (liu2022neroic) or from distributed agents in a federated learning setting .
+The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems .
+Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis or respiratory healthcare sensing , is key.
 3. Methodology
@@
 The networks are trained end-to-end by minimizing a composite loss function.
-The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production and WiFi-based sensing (wang2024targetoriented).
+The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production and WiFi-based sensing .
 3.4.1. Photometric Reconstruction Loss
@@
 We also use a mask loss to enforce that SDF values inside the object (as defined by the input masks) are negative.
-This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains (huang2024phyfinatt).
+This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains .
 3.4.3. Material and Lighting Priors
@@
 We also regularize the total intensity and smoothness of the predicted environment light map from $f_{light}$ to prevent overly complex lighting from being ”baked” into the material albedo.
-The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning (liu2024finch) or visual understanding .
+The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning or visual understanding .
 3.4.4. Total Loss
@@
 We evaluate our method on both synthetic and real-world datasets to demonstrate its effectiveness and generalization capabilities.
-The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language to ensuring robust performance in WiFi-based activity recognition under interference (huang2023phaseanti).
+The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language to ensuring robust performance in WiFi-based activity recognition under interference.
 Synthetic Dataset
@@
 Since ground-truth material properties are not available for these real objects, evaluation is primarily focused on the quality of the geometric reconstruction and the visual fidelity of novel view synthesis and relighting results.
-The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals (gu2023attentionbased).
+The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals .
 4.1.2. Evaluation Metrics
@@
   • Novel View Synthesis: For both datasets, we evaluate the quality of rendered novel views against ground-truth images using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and the Learned Perceptual Image Patch Similarity (LPIPS) (srinivasan2023nero).
-PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog or emotion recognition (gu2023wife).
+PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog or emotion recognition .
   • Material Map Accuracy: On the synthetic dataset, where ground-truth PBR maps are available, we compute the PSNR between our decomposed material maps (albedo, roughness, metallic) and the ground-truth maps.
@@
 The loss weights are set empirically, with $\lambda_{L1}=1.0$, $\lambda_{LPIPS}=0.2$, $\lambda_{eikonal}=0.1$, and $\lambda_{mat}=0.01$.
-The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning (liu2024finch; liu2024federated) or dealing with the security implications of advanced sensing technologies (huang2024keystrokesniffer).
+The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning or dealing with the security implications of advanced sensing technologies .
 Our code and pre-trained models will be made publicly available to facilitate future research.
@@
 This is in contrast to geometry-only methods like NeuS, whose view-dependent appearance model can sometimes produce results that look flat or lack realism under complex lighting.
-The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks (liu2024federated) to secure WiFi-based gesture recognition (gu2023attentionbased).
+The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks
 Table 2. Quantitative comparison for novel view synthesis on the DTU dataset. We report PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$). Best results are in bold.
@@
 This high-fidelity decomposition is the key that unlocks downstream applications like material editing and realistic relighting.
-The clarity of our decomposition can be likened to the goal of achieving interference-independent feature extraction in WiFi sensing (huang2023phaseanti).
+The clarity of our decomposition can be likened to the goal of achieving interference-independent feature extraction in WiFi sensing.
 Figure 2.
@@
 This raises important questions about privacy and security.
-The ability to reconstruct detailed 3D models could be misused, and the underlying sensing systems could be vulnerable, a concern echoed in the domain of WiFi sensing which has demonstrated risks like keystroke sniffing (huang2024keystrokesniffer).
-Future work should explore privacy-preserving reconstruction techniques, potentially leveraging federated learning frameworks (liu2024federated) to train models without centralizing sensitive user data, and ensuring the security of the underlying capture and processing pipeline.
+The ability to reconstruct detailed 3D models could be misused, and the underlying sensing systems could be vulnerable, a concern echoed in the domain of WiFi sensing which has demonstrated risks like keystroke sniffing .
+Future work should explore privacy-preserving reconstruction techniques, potentially leveraging federated learning frameworks to train models without centralizing sensitive user data, and ensuring the security of the underlying capture and processing pipeline.